What is the flow rate for 50 ml of an antibiotic IVPB infused over 30 minutes if the administration set is calibrated to deliver 15 gtt/mL? Use the following information for questions 21 and 22 A physician orders 120 units of insulin in 1,000 mL of DsW to be administered over S hr 20. How many drops per minute should be administered if the set is calibrated to deliver 40 gtt/mli? 21 How much insulin will be delivered in the first 75 min of the infusioni? Use the following information for problems 23 and 24. A 112-ib nursing home patient is presecribed 1o mg/ng of phenrytoin administered at a ate of 50 mg/min. The technician % NS. prepares an IV with a concentration of 1 g phenytoin in 1 L of 0.9 22. How long would this IV last? 23. If the dose was increased to 15 mg/kg, what would the IV flow rate be in ml/hr? Use the following information for problems 25 and 26 The doctor orders 4 mg/kg/hr of Phenobarbital IV for an 88-lb patient who is having a seizure. The pharmacist instructs the technician to put 100 mg of Phenobarbital in 500 mL of DsW. 24. What would the IV flow rate be in mL/hr for this preparation? 25. How many minutes would one IV last? If the final volume of amoxicillin reconstituted suspension should be 150 mL, and the instructions indicate to add 134 mL of water to the dry powder, what is the powder volume of the amoxicillin?
Answer #1

ANSWER:

19.

Antibiotic to be infused in 30 min=50 ml

Antibiotic to be infused in 1 min =50/30=1.6ml

Flow rate is 15drops in 1ml

so, 1.6ml =15*1.6

=24 drops.

so, flow rate is 24drops per minute.

20.

1000ml to be infused in 8*60=480 min.

1ml =40 drops (given)

so, 1000ml =40*1000

=40,000 drops.

so, in 480 min 40,000 drops to be adiminsterd.

thus in 1 min =40,000/480=83 drops

Hence, 83 drops per minute to be administerd to the patient.

21.

8 hrs= 8*60min

=480 min.

In 480min the Insulin solution Administered is =1000ml

In 75 Min the insulin Solution administered =1000/480*75

=156.25 or 156 ml approx.

thus in 1000ml there is 126 units of insulin

so, in 156 ml there is 126/1000*156

= 19.6 units

so, 19.6 units of insulin is deliveredin first 75 min of infusion.